OpenStudy (y1066380):

how to solve d/dx(1-sinx)?

OpenStudy (anonymous):

\[ \frac{d}{dx}(1-\sin(x)) = \frac{d}{dx}(1) - \frac{d}{dx}(\sin (x)) \]

OpenStudy (gorv):

\[\frac{ d (1-sinx)}{ dx }=\frac{ d (1) }{ dx } - \frac{ d sinx }{ dx }=0-cosx\]